OpenStudy (anonymous):

how long should an escalator be if it is to make an angle of 33 degrees with the floor and carry people a vertical distance of 21 feet between floors?

OpenStudy (anonymous):

 would this pic be right ?

OpenStudy (anonymous):

yes thats right, you know the opposite side and you want to find the hypotenuse, so which of sin cos tan do you need to use?

OpenStudy (anonymous):

38 :) use formula sine33= opp21/hyp

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

wouldnt it be hypotenuse/opposite since we dont know the hypotenuse

OpenStudy (anonymous):

hypotenuse/opposite gives us 1/sin not sin,

OpenStudy (anonymous):

not unless your doing scs or whatever is the opposite of sin lol. sin=opp/hyp

OpenStudy (anonymous):

38.55 is the hyp. so that would actually be 39. if that's what youre looking for

OpenStudy (anonymous):

so how do i write that out and solve then \[\sin 33=21/h\] so how do i solve for this

OpenStudy (anonymous):

@ninamarie

OpenStudy (anonymous):

@JamesWolf

OpenStudy (anonymous):

sin 33 = 21/h h (sin 33) = 21 h = 21/sin 33

OpenStudy (anonymous):

did you get the answer or do you ant it explained differently?

OpenStudy (anonymous):

want*

OpenStudy (anonymous):

i got it thanks :D

OpenStudy (anonymous):

okay :)